OpenStudy (anonymous):

Solve the equation 2cos^2x = sin2x , giving your answer in terms of π. 0

OpenStudy (anonymous):

use double angle formula for the right side to get: 2cos^2x = 2sinxcosx 2cos^2x - 2sinxcosx = 0 2cosx(cosx - sinx) = 0 so cosx = 0 gives pi/2. cosx - sinx = 0 or tanx = 1 gives pi/4.

OpenStudy (callisto):

can i try this way? 2cos^2x = sin2x cos2x +1 = sin2x cos2x - sin2x =-1 (cos2x - sin2x)^1 = (-1)^2 (cos2x)^2 -2cos2xsin2x + (sin2x)^2 =1 1-2sin2xcos2x =1 -sin4x = 0 sin4x =0 4x = π x = π/4

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

np